# FedNoisy: Federated Noisy Label Learning Benchmark

### v1

@@ 4.2 Datasets and Data Heterogeneity Simulation
As summarized in Table 1, FedNoisy serves 5 datasets for variant tasks. MNIST and SVHN are digit recognition tasks where digits 1 to 10 are included. CIFAR-10 and CIFAR-100 are popular object recognition datasets with 10 and 100 classes, respectively. These four datasets are frequently adopted in both FL and CNLL areas. We apply globalized and localized noise scenes to them and evaluate them on the clean test set. Clothing1M is a large-scale clothing image dataset of 14 categories from shopping websites, containing natural label noise at a ratio of 39.46%. It is a popular dataset widely used in CNLL, and we use it for the real-world noise scene in FedNoisy. Following DivideMix [31], we randomly sample a class-balanced subset containing 64k images from the original noisy training set as the noisy global dataset $\tilde{\mathcal{D}}$ for training, and report the test performance on 10k clean test set.
Table 1: Summary of datasets provided in FedNoisy
@@
Globalized/Localized MNIST [53] 60K N/A 10K 10 $28\times 28$ N/A
SVHN [54] 73K N/A 26K 10 $32\times 32\times 3$ N/A
CIFAR-10 [55] 50K N/A 10K 10 $32\times 32\times 3$ N/A
CIFAR-100 [55] 50K N/A 10K 100 $32\times 32\times 3$ N/A
Real-world Clothing1M [51] 1.0M 14K 10K 14 $224\times 224\times 3$ $\approx 39.46$
Found by massive existing FL research [50, 56, 46], data heterogeneity is a common yet especially challenging scenario in practice where each client has non-identical-and-independent (non-IID) data distributions compared to others. Given a data sample $(x,y)$, non-IID indicates the joint probability distribution $\mathcal{P}(x,y)$ is different for different client $i$ and $j$ ($i,j\in[K]$, $i\neq j$), that is, $\mathcal{P}_{i}(x,y)\neq\mathcal{P}_{j}(x,y)$. In FedNoisy, we enclose three representative non-IID schemes together with the standard IID setting, as follows. 1) Quantity skew. Different clients hold different amounts of data, controlled by Dirichlet distribution with parameter $\alpha$, denoted as noniid-quantity=$\alpha$. 2) Dirichlet label skew. Different clients hold local datasets with different proportions of different classes, controlled by Dirichlet distribution with parameter $\alpha$. We denote this as noniid-labeldir=$\alpha$. 3) Label-quantity skew. In this Non-IID case, the class heterogeneity is simulated by assigning samples of a specified number of classes to each client. We use noniid-#label=c to denote this partition given class number c. Detailed descriptions of the data partition are presented in Appendix B.
4.3 Implemented methods
Table 2: Summary of baselines in FedNoisy
@@ 4.3 Implemented methods
Robust regularization Mixup [36]
Robust loss function SCE [37]
GCE [57]
MAE [58]
Loss adjustment M-DYR-H [40]
M-DYR-S [40]
DM-DYR-SH [40]
Sample selection Co-teaching [41]
In FedNoisy, we integrate representative NLL algorithms in the local client training, which can work seamlessly with most federated learning algorithms. The list of algorithms is summarized in Table 2. Previous NLL methods can be roughly grouped into four categories: i) robust regularization methods, ii) robust loss function design, iii) loss adjustment methods, and iv) sample selection methods. Though many algorithms have been developed under these principles, they are not extensively used as baselines in existing FNLL research. The main reason could be a lack of a well-developed codebase, leading to formidable burdens including them for comparisons. In our benchmark, we tackle the critical challenge by providing a package of ready-to-go codes for these algorithms. Importantly, we implement these algorithms mainly on the client side with least modifications on the global communication. The decoupled code structure makes these algorithms easy to be extended with more advanced federated aggregations.
5 Experiments
@@
5.1 Benchmark performance
This section presents the performance for 20 benchmark settings in FedNoisy. We compare the test accuracy of 20 benchmark settings in FedNoisy as an overview of the challenges in FNLL. In the experiment, we select some default hyper-parameters: $\varepsilon_{global}=0.4$ for globalized noise. For localized noise, we use $\varepsilon_{k}\sim\mathcal{U}(0.3,0.5)$ (that is, $\varepsilon_{local}=0.4$, $\sigma=0.1$). The parameters of localized noise yield a fair comparison between localized and globalized settings with similar overall noisy samples, which is shown in Appendix C). For data partition, we choose iid, noniid-#label=3, noniid-labeldir=0.1 and noniid-quantity=0.1 as the default settings. All analyses are based on the results of CIFAR-10 on FedAvg of 10 clients using VGG-16 [59] with 500 communication rounds and 5 local epochs. We report the averaged test accuracy over the last 10 rounds, and each experiment is repeated 3 times with different random seeds. More details on experiments can be found in Appendix A. Table 3 presents the complete results for 20 default settings. Results for the same experiment setting for more datasets can be found in Appendix E.
Table 3: Test accuracy of CIAFR-10 on 10 clients using FedAvg with VGG16. Significant gaps exist between different noise schemes, especially under heterogeneous data partitions.
@@ 5.1 Benchmark performance
5.2 Fine-grained Ablation Studies of FNLL
@@
For globalized noise, we choose $\varepsilon_{global}$ from $\{0.1,0.2,0.3,0.4,0.5,0.6,0.7\}$ for symmetric noise and $\varepsilon_{global}$ from $\{0.1,0.2,0.3,0.4,0.5,0.6\}$ for asymmetric noise, while keeping other learning setting same as default settings. For localized noise, choose $\varepsilon_{local}$ from $\{0.1,0.2,0.3,0.4,0.5\}$ and $\sigma=0.1$ for both symmetric and asymmetric noise. And to make sure localized settings with large noise ratios converge, we train FedAvg for 700 rounds while keeping other learning settings default. Figure 3 shows the curves of test accuracy versus noise ratio for both globalized and localized noise settings. Complete results are in Appendix F.1.
Figure 2: Accuracy for different noise ratios on CIFAR-10 with 10 clients. The $x$-axis noise ratio is $\varepsilon_{global}$ for globalized noise, and $\varepsilon_{local}$ for localized noise $\mathcal{U}(\varepsilon_{local}-0.1,\varepsilon_{local}+0.1)$. From left to right: globalized noise, localized noise.
Figure 2: Accuracy for different noise ratios on CIFAR-10 with 10 clients. The $x$-axis noise ratio is $\varepsilon_{global}$ for globalized noise, and $\varepsilon_{local}$ for localized noise $\mathcal{U}(\varepsilon_{local}-0.1,\varepsilon_{local}+0.1)$. From left to right: globalized noise, localized noise.
Figure 3: Sensitivity on different noise ratios on CIFAR-10 with 10 clients. The $x$-axis noise ratio is $\varepsilon_{global}$ for globalized noise, and $\varepsilon_{local}$ for localized noise $\mathcal{U}(\varepsilon_{local}-0.1,\varepsilon_{local}+0.1)$. From left to right: globalized noise, localized noise.
1) Symmetric noise is harder than asymmetric with Non-IID. In the centralized setting, it has been found that asymmetric label noise is harder than symmetric noise, especially when the noise ratio is larger than $0.5$, considering each class has more correct samples to learn from under symmetric noise [41, 31]. However, in FNLL, the conclusion depends on the partition type. As shown in Figure 3, in both globalized and localized settings, performance with symmetric noise tends to become better than that with asymmetric noise when the noise ratio is larger than $0.2$ under IID partition (class-balanced on each client in this experiment), and the performance gap increases as noise ratio grows. While in Non-IID which class-imbalance exists (i.e., noniid-labeldir and noniid-#label), symmetric noise is more difficult than asymmetric noise even with a small noise ratio. And this performance gap also becomes larger as the noise ratio grows. This intriguing phenomenon might be attributed to the class imbalance problem in these label distribution skew scenes, which encourages researchers to alleviate class imbalance when dealing with symmetric noise in FL settings.
@@
Figure 4: Accuracy drop ratio from IID to Non-IID with globalized noise on CIFAR-10. Give noise ratio $\varepsilon$, the accuracy drop ratio of certain non-iid partition at $\varepsilon$ is calculated by $[\mathrm{acc}_{\texttt{iid}}(\varepsilon)-\mathrm{acc}_{\texttt{non-iid}}(\varepsilon)]/\mathrm{acc}_{\texttt{iid}}(\varepsilon)$. The dotted lines mark the values when $\varepsilon=0.0$, and the dash-dotted lines mark accuracy drop ratio equals 0. $\varepsilon$ is $\varepsilon_{global}$ for globalized noise, and $\varepsilon_{local}$ for localized noise $\mathcal{U}(\varepsilon_{local}-0.1,\varepsilon_{local}+0.1)$. From left to right: globalized noise, localized noise.
@@
$s(\varepsilon)=\frac{\mathrm{acc}(\varepsilon)-\mathrm{acc}(\varepsilon+\Delta)}{\Delta},$ (4)
@@
4) Label noise increases the gradient norm in FL. Previous CNLL works reveal that noisy labels can influence characteristics of the training process: [41, 31] believe that noisy samples tend to have larger loss values compared with clean ones; [60, 61] showed that noisy samples increase gradient norm of the training batch. Some FNLL papers have utilized sample loss for noisy sample detection in FNLL, while none has explored noisy labels’ impact on gradient norm in FNLL yet. We calculate the per-round global model gradient norm as $\|w^{t}-w^{t-1}\|_{2}$ for each round and perform FedAvg training for iid, noniid-#label, and noniid-labeldir with clean or globalized/localized noise. We use a learning rate of $0.005$ for all different settings for fair comparison in this experiment. Figure 5 shows the results of globalized settings, and Figure 7 in Appendix D shows the results of localized settings. We conclude from these two figures: 1) the gradient norm converges during the training process; 2) noisy labels enlarge the gradient norm of the global model under both IID and Non-IID partitions; 3) given the same noise setting, the curve of gradient norm under IID fluctuates significantly compared with Non-IID partitions. From these observations, we conjecture that the gradient norm is also ponderable information that can be leveraged in FNLL.
Figure 5: Gradient norm of global model for clean and globalized noise $\varepsilon_{global}=0.4$ settings.
6 Conclusion
In this paper, we propose comprehensive data settings and standardized simulation pipeline for FNLL in benchmark FedNoisy, including 20 basic data settings for potential FNLL settings, 5 datasets, and 8 baselines covering a wide scope of methodologies. We conduct extensive comparisons and reveal some special characteristics of FNLL, which we hope can motivate further research on noisy label learning in FL community. Currently, FedNoisy is implemented on PyTorch [62] and FedLab [45], which might limit the usage of researchers familiar with other ML backends and FL frameworks, such as TensorFlow [63], Flower [64], FedML [65]. Supporting other backends/frameworks is one of the future directions. We will keep updating FedNoisy to support more datasets and FNLL baselines.
@@ Appendix A Experiment details
We use VGG-16 [59] as the base model for CIFAR-10 and SVHN experiments, and use CNN for MNIST following [46]. We use SGD optimizer with momentum $0.9$, default weight decay as $5\mathrm{e}{-4}$, and the learning rate is tuned from $\{0.01,0.005,0.002,0.001\}$ for all settings. The batch size is set to 128 and the number of local epochs is set to 5 by default. For basic 20 settings (clean, $\varepsilon_{global}=0.4$ and $\mathcal{U}(0.3,0.5)$ with 4 partitions) on all datasets, we use 500 rounds training. For experiments with different noise ratios on CIFAR-10, we use 500 rounds for globalized noise, and 700 rounds for localized noise for convergence considerations.
Our FNLL simulation pipelines are built upon FedLab [45] framework and PyTorch [62]. And we release the codes with Apache License 2.0.

### v4

@@ 4.2. Datasets and Data Heterogeneity Simulation
As summarized in Table 1, FedNoisy serves 5 datasets for variant tasks. MNIST and SVHN are digit recognition tasks where digits 1 to 10 are included. CIFAR-10 and CIFAR-100 are popular object recognition datasets with 10 and 100 classes, respectively. These four datasets are frequently adopted in both FL and CNLL areas. We apply globalized and localized noise scenes to them and evaluate them on the clean test set. Clothing1M is a large-scale clothing image dataset of 14 categories from shopping websites, containing natural label noise at a ratio of 39.46%. It is a popular dataset widely used in CNLL, and we use it for the real-world noise scene in FedNoisy. Following DivideMix (Li et al., 2020), we randomly sample a class-balanced subset containing 64k images from the original noisy training set as the noisy global dataset $\tilde{\mathcal{D}}$ for training, and report the test performance on 10k clean test set.
Table 1. Summary of datasets provided in FedNoisy
@@
Globalized/Localized MNIST (LeCun, 1998) 60K N/A 10K 10 $28\times 28$ N/A
SVHN (Netzer et al., 2011) 73K N/A 26K 10 $32\times 32\times 3$ N/A
CIFAR-10 (Krizhevsky et al., 2009) 50K N/A 10K 10 $32\times 32\times 3$ N/A
CIFAR-100 (Krizhevsky et al., 2009) 50K N/A 10K 100 $32\times 32\times 3$ N/A
Real-world Clothing1M (Xiao et al., 2015) 1.0M 14K 10K 14 $224\times 224\times 3$ $\approx 39.46$
WebVision (Li et al., 2017) 2.4M 50K 50K 1000 $256\times 256\times 3$ $\approx 20.00$
Found by massive existing FL research (McMahan et al., 2017; Kairouz et al., 2021; Li et al., 2022a), data heterogeneity is a common yet especially challenging scenario in practice where each client has non-identical-and-independent (non-IID) data distributions compared to others. Given a data sample $(x,y)$, non-IID indicates the joint probability distribution $\mathcal{P}(x,y)$ is different for different client $i$ and $j$ ($i,j\in[K]$, $i\neq j$), that is, $\mathcal{P}_{i}(x,y)\neq\mathcal{P}_{j}(x,y)$. In FedNoisy, we enclose three representative non-IID schemes together with the standard IID setting, as follows. 1) Quantity skew. Different clients hold different amounts of data, controlled by Dirichlet distribution with parameter $\alpha$, denoted as noniid-quantity=$\alpha$. 2) Dirichlet label skew. Different clients hold local datasets with different proportions of different classes, controlled by Dirichlet distribution with parameter $\alpha$. We denote this as noniid-labeldir=$\alpha$. 3) Label-quantity skew. In this Non-IID case, the class heterogeneity is simulated by assigning samples of a specified number of classes to each client. We use noniid-#label=c to denote this partition given class number c. Detailed descriptions of the data partition are presented in Appendix B.
4.3. Implemented methods
Table 2. Summary of baselines in FedNoisy
@@ 4.3. Implemented methods
Robust regularization Mixup (Zhang et al., 2018)
Robust loss function SCE (Wang et al., 2019)
GCE (Zhang and Sabuncu, 2018)
MAE (Ghosh et al., 2017)
Loss adjustment M-DYR-H (Arazo et al., 2019)
M-DYR-S (Arazo et al., 2019)
DM-DYR-SH (Arazo et al., 2019)
Sample selection Co-teaching (Han et al., 2018)
FNLL FedNoRo (Wu et al., 2023)
In FedNoisy, we integrate representative NLL algorithms in the local client training, which can work seamlessly with most federated learning algorithms. The list of algorithms is summarized in 2. Previous NLL methods can be roughly grouped into four categories: i) robust regularization methods, ii) robust loss function design, iii) loss adjustment methods, and iv) sample selection methods. Though many algorithms have been developed under these principles, they are not extensively used as baselines in existing FNLL research. The main reason could be a lack of a well-developed codebase, leading to formidable burdens including them for comparisons. In our benchmark, we tackle the critical challenge by providing a package of ready-to-go codes for these algorithms. Importantly, we implement these algorithms mainly on the client side with the least modifications on the global communication. The decoupled code structure makes these algorithms easy to be extended with more advanced federated aggregations.
5. Experiments
@@
5.1. Benchmark performance
This section presents the performance for 20 benchmark settings in FedNoisy. We compare the test accuracy of 20 benchmark settings in FedNoisy as an overview of the challenges in FNLL. In the experiment, we select some default hyper-parameters: $\varepsilon_{global}=0.4$ for globalized noise. For localized noise, we use $\varepsilon_{k}\sim\mathcal{U}(0.3,0.5)$ (that is, $\varepsilon_{local}=0.4$, $\sigma=0.1$). The parameters of localized noise yield a fair comparison between localized and globalized settings with similar overall noisy samples, which is shown in Appendix C). For data partition, we choose iid, noniid-#label=3, noniid-labeldir=0.1 and noniid-quantity=0.1 as the default settings. All analyses are based on the results of CIFAR-10 on FedAvg of 10 clients using VGG-16 (Simonyan and Zisserman, 2015) with 500 communication rounds and 5 local epochs. We report the averaged test accuracy over the last 10 rounds, and each experiment is repeated 3 times with different random seeds. More details on experiments can be found in Appendix A. Table 3 presents the complete results for 20 default settings. Results for the same experiment setting for more datasets can be found in Appendix E.
Table 3. Test accuracy of CIAFR-10 on 10 clients using FedAvg with VGG16. Significant gaps exist between different noise schemes, especially under heterogeneous data partitions.
@@ 5.1. Benchmark performance
5.2. Baseline Performance
We evaluate all baselines using MNIST, CIFAR-10, and Clothing1M under noniid-#label setting. We aim to explore the trade-off between local computation cost and model performance. The scatter visualizations for CIFAR-10 and Clothing1M are shown in Figure 2 and Figure 3, and the result for MNIST is shown in Figure 11 of Appendix G for space limit. We categorize the baselines based on their ranking into tier 1 (T1), tier 2 (T2), and others for accuracy and computation time, respectively. We use red dashed boxes for the accuracy tiers and green dashed boxes for time.
Generalizability over different model architectures.
For CIFAR-10, we find that the MAE yields reasonable performance with an 8-layer CNN, but fails to converge on VGG16. This inconsistency underscores the need for researchers to consider the generalizability of the FNLL across various model architectures.
Baseline robustness across datasets and noise settings.
By comparing Figure 11, Figure 2, and Figure 3, we find that the performance rankings of baselines exhibit variability across different datasets as well as noise settings, indicating a lack of robustness. For example, M-DYR-H is in accuracy T2 for globalized noise, while only achieving only comparable with naive baseline under localized noise on CIFAR-10, as shown in Figure 2. FedNoRo is in accuracy T1 on Clothing1M as shown in Figure 3, while is even worse than the vanilla baseline for both CIFAR-10 and MNIST. This indicates that researchers should pay more attention to the algorithms’ robustness over datasets and noise settings when working on FNLL.
Accuracy and computation trade-off.
Based on results in Figure 11, Figure 2, and Figure 3, we assign points to baselines based on their accuracy/computation tier: 1 point for tier 1 (T1) reflecting either accuracy or local computation time, 2 points for tier 2 (T2), and 4 points for the remaining. We aggregate these points across all datasets to derive a final score for each baseline, assessing the trade-off between accuracy and computational efficiency. A lower score indicates superior overall accuracy or lower overall computational costs. As shown in Table 4, SCE and GCE achieve the best trade-off between accuracy and computation cost  [^2] We exclude MAE as the best baseline due to its limited generalizability across different model architectures.. Consequently, we recommend that researchers consider both accuracy and computational cost, and adopt SCE and GCE as strong baselines for both criteria when developing FNLL algorithms.
[…]
Vanilla 18 5
Mixup 12 5
SCE 7 5
GCE 9 5
MAE 8 5
M-DYR-H 10 14
M-DYR-S 12 14
M-DYR-SH 12 14
Co-teaching 8 12
FedNoRo 17 5
Table 4. Accuracy and time trade-off scores for different baselines over datasets MNIST, CIFAR-10, and Clothing1M. The blue background indicates the baselines with the best trade-off between accuracy and computation cost. indicates the baseline has trouble with convergence when using different model architectures.
Figure 2. The trade-off between accuracy and per-round local computation time on CIFAR-10 on 10 clients under noniid-#label=3 partition with symmetric noise. From left to right: globalized noise, localized noise. Baselines with * are using 8-layer CNN following (Ma et al., 2020), otherwise with default VGG16.
Figure 3. The trade-off between accuracy and per-round local computation time on Clothing1M on 10 clients under noniid-#label=5 partition with real noise. Red and green dash boxes are used to categorize the baselines into several tiers for both test accuracy and per-round time, respectively. T1 denotes tier 1, and T2 denotes tier 2.
5.3. Fine-grained Ablation Studies of FNLL
Figure 4. Gradient norm of global model for clean and globalized noise $\varepsilon_{global}=0.4$ settings.
@@ 5.3. Fine-grained Ablation Studies of FNLL
For globalized noise, we choose the global noise ratio $\varepsilon_{global}$ from $\{0.1,0.2,0.3,0.4,0.5,0.6,0.7\}$ for symmetric noise, while $\{0.1,0.2,0.3,0.4,0.5,0.6\}$ for asymmetric noise, while keeping other learning setting same as default settings. For localized noise, choose $\varepsilon_{local}$ from $\{0.1,0.2,0.3,0.4,0.5\}$ and $\sigma=0.1$ for both symmetric and asymmetric noise. And to make sure localized settings with large noise ratios converge, we train FedAvg for 700 rounds while keeping other learning settings default. Figure 6 shows the curves of test accuracy versus noise ratio for both globalized and localized noise settings. Complete results are in Appendix F.1.
Figure 5. Accuracy for different noise ratios on CIFAR-10 with 10 clients. The $x$-axis noise ratio is $\varepsilon_{global}$ for globalized noise, and $\varepsilon_{local}$ for localized noise $\mathcal{U}(\varepsilon_{local}-0.1,\varepsilon_{local}+0.1)$. From left to right: globalized noise, localized noise.
Figure 6. Sensitivity on different noise ratios on CIFAR-10 with 10 clients. The $x$-axis noise ratio is $\varepsilon_{global}$ for globalized noise, and $\varepsilon_{local}$ for localized noise $\mathcal{U}(\varepsilon_{local}-0.1,\varepsilon_{local}+0.1)$. From left to right: globalized noise, localized noise.
1) Symmetric noise is harder than asymmetric with Non-IID. In the centralized setting, it has been found that asymmetric label noise is harder than symmetric noise, especially when the noise ratio is larger than $0.5$, considering each class has more correct samples to learn from under symmetric noise (Han et al., 2018; Li et al., 2020). However, in FNLL, the conclusion depends on the partition type. As shown in Figure 6, in both globalized and localized settings, performance with symmetric noise tends to become better than that with asymmetric noise when the noise ratio is larger than $0.2$ under IID partition (class-balanced on each client in this experiment), and the performance gap increases as noise ratio grows. While in Non-IID which class-imbalance exists (i.e., noniid-labeldir and noniid-#label), symmetric noise is more difficult than asymmetric noise even with a small noise ratio. And this performance gap also becomes larger as the noise ratio grows. This intriguing phenomenon might be attributed to the class imbalance problem in these label distribution skew scenes, which encourages researchers to alleviate class imbalance when dealing with symmetric noise in FL settings.
@@
Figure 7. Accuracy drop ratio from IID to Non-IID with globalized noise on CIFAR-10. Give noise ratio $\varepsilon$, the accuracy drop ratio of certain non-iid partition at $\varepsilon$ is calculated by $[\mathrm{acc}_{\texttt{iid}}(\varepsilon)-\mathrm{acc}_{\texttt{non-iid}}(\varepsilon)]/\mathrm{acc}_{\texttt{iid}}(\varepsilon)$. The dotted lines mark the values when $\varepsilon=0.0$, and the dash-dotted lines mark accuracy drop ratio equals 0. $\varepsilon$ is $\varepsilon_{global}$ for globalized noise, and $\varepsilon_{local}$ for localized noise $\mathcal{U}(\varepsilon_{local}-0.1,\varepsilon_{local}+0.1)$. From left to right: globalized noise, localized noise.
@@
(4) $s(\varepsilon)=\left[\mathrm{acc}(\varepsilon)-\mathrm{acc}(\varepsilon+\Delta)\right]/\Delta,$
@@
4) Label noise increases the gradient norm in FL. Previous CNLL works reveal that noisy labels can influence characteristics of the training process: (Han et al., 2018; Li et al., 2020) believe that noisy samples tend to have larger loss values compared with clean ones; (Liu et al., 2021b; Wei et al., 2022) showed that noisy samples increase gradient norm of the training batch. Some FNLL papers have utilized sample loss for noisy sample detection in FNLL, while none has explored noisy labels’ impact on gradient norm in FNLL yet. We calculate the per-round global model gradient norm as $\|w^{t}-w^{t-1}\|_{2}$ for each round and perform FedAvg training for iid, noniid-#label, and noniid-labeldir with clean or globalized/localized noise. We use a learning rate of $0.005$ for all different settings for fair comparison in this experiment. Figure 4 shows the results of globalized settings, and Figure 9 in Appendix D shows the results of localized settings. We conclude from these two figures: 1) the gradient norm converges during the training process; 2) noisy labels enlarge the gradient norm of the global model under both IID and Non-IID partitions; 3) given the same noise setting, the curve of gradient norm under IID fluctuates significantly compared with Non-IID partitions. From these observations, we conjecture that the gradient norm is also ponderable information that can be leveraged in FNLL.
@@ Appendix A Experiment details
We use VGG-16 (Simonyan and Zisserman, 2015) as the base model for CIFAR-10 and SVHN experiments, and use CNN for MNIST following (Li et al., 2022a). We use SGD optimizer with momentum $0.9$, default weight decay as $5\mathrm{e}{-4}$, and the learning rate is tuned from $\{0.01,0.005,0.002,0.001\}$ for all settings. The batch size is set to 128 and the number of local epochs is set to 5 by default. For basic 20 settings (clean, $\varepsilon_{global}=0.4$ and $\mathcal{U}(0.3,0.5)$ with 4 partitions) on all datasets, we use 500 rounds training. For experiments with different noise ratios on CIFAR-10, we use 500 rounds for globalized noise, and 700 rounds for localized noise for convergence considerations.
Our FNLL simulation pipelines are built upon FedLab (Zeng et al., 2023) framework and PyTorch (Paszke et al., 2019). And we release the codes with Apache License 2.0.